# From the Stochastic Embedding Sufficiency Theorem to a Superspace Diffusion Framework

### 20 Mar 2026 v1

@@ 3.6 Quantum Field Theory: The Roles of $\hbar$, $c$, and $G$
The pipeline was applied at seven frequencies spanning a decade ($\omega\in[10^{12},10^{13}]$ rad/s) at fixed damping $\gamma=10^{11}$ s-1. The recovered $\hat{\sigma}$ was constant across all frequencies to within 1% (Fig. 9c), with mean error $0.50\%$ and maximum error $1.00\%$ from 20 independent random seeds (Fig. 9d).
@@
For the photon, $\gamma$ is an external parameter (cavity losses, material absorption). For gravity, however, the coupling is not external: the gravitational field is self-interacting (metric fluctuation modes carry energy, and that energy gravitates by the equivalence principle), so the self-coupling rate is uniquely determined by the remaining constant, Newton’s gravitational constant $G$, through $\gamma_{\mathrm{grav}}=1/t_{P}$, where $t_{P}=\sqrt{\hbar G/c^{5}}$ is the Planck time [26]. Substitution yields $\sigma=c\sqrt{1/t_{P}}=\ell_{P}/\sqrt{t_{P}}$, which corresponds to a proper-distance fluctuation of one Planck length per Planck time. This inference is developed in the Discussion (§4) and formalised in Appendix C.
@@ 3.7 Summary of Quantitative Results
Table 2 consolidates the principal recovered quantities, errors, and statistical assessments for all nine domains.
Table 2: Summary of quantitative results. Recovery errors range from $0.026\%$ to $\sim\!1\%$. All $p$-values exceed $0.05$: no null hypothesis is rejected. Ensemble CI coverage is reported for stochastic free-run ensembles (250 realisations), where values near 95% indicate well-calibrated uncertainty. Tests are ordered by the progression of complexity described in §2.4.
Test Domain Theoretical $\sigma$ Error pp-value 95% CI coverage
C1 Classical mechanics $\sigma=0$ $10^{-13}$ 0.92 (deterministic)
C4 Quantum mechanics $\sigma\approx 0$; $\hbar$ in $\mu$ 0.026% 0.84 (deterministic)
C6 Electromagnetism $\sigma=0$ $10^{-13}$ 0.91 (deterministic)
C2 Statistical mechanics $\sqrt{2\gamma k_{B}T/m}$ 0.79% 0.34 94.5%
C3 Nuclear physics $\sqrt{\mu}$ (Fano =1=1) 0.08% 0.50 95.8%
C8 Quantum field theory $\sqrt{\gamma\hbar\omega/m}$ 0.35%
C5 Chemical kinetics $\propto\Omega^{-1/2}$ 0.38% 0.46 94.0%
C7 Relativistic QM $\sigma_{0}/\gamma$ 0.66% 0.58
C9 QED (photon field) $c\sqrt{\gamma}$ 0.50%
@@ 4.1 The $\sigma$-Continuum as Empirical Evidence for Universal Stochastic Structure
The nine-domain validation of §3 establishes that the stochastic differential equation $\mathrm{d}X=\mu(X)\,\mathrm{d}t+\sigma(X)\,\mathrm{d}W$ is not solely a mathematical device but a recurring structural element across physical domains. A single non-parametric pipeline, receiving only raw time series and sampling interval, recovers the governing equations of classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, and quantum field theory, each to sub-percent accuracy and each without domain-specific input.
@@
The $\sigma$-continuum (Fig. 1) is an observation arising from these nine independent experiments. The natural next step is to determine whether the pattern extends to gravitational degrees of freedom, and if so, what determines $\sigma$ for gravity.
@@ 4.2 Relativistic Structure in the Diffusion Coefficient
The pipeline recovers the $1/\gamma$ suppression to 0.66% (Fig. 8d) and the group velocity to $R^{2}=0.9999$ (Fig. 8f). The independent recovery of $v_{g}$ at each momentum mitigates a potential circularity: because $\gamma$ depends on $v$, the suppression $\sigma\propto 1/\gamma$ is validated only if the velocity is itself recovered independently, not assumed. Both channels, drift and diffusion, carry relativistic content. The speed of light $c$ enters $\sigma$ through the dispersion relation, alongside $\hbar$ and $m$.
Five structural observations follow from C6–C7 taken together.
@@
Fifth, the progression from C1 ($\sigma=0$) through C2–C5 ($\sigma>0$, non-relativistic) to C7 ($\sigma=\sigma_{0}/\gamma$, relativistic) constitutes an empirically observed hierarchy. The subsequent extrapolation to gravitational degrees of freedom (§4.3) is therefore a continuation of this hierarchy, not a dimensional-analysis leap.
@@ 4.3 From Quantum Fields to Gravitational $\sigma$
where $t_{P}=\sqrt{\hbar G/c^{5}}$ is the Planck time. Substitution into $\sigma=c\sqrt{\gamma}$ gives
@@
The derivation thus proceeds in three stages, each building on validated results. Stage 1: the quantum harmonic oscillator (C8) establishes the fluctuation–dissipation relation $\sigma^{2}=\gamma\hbar\omega/m$ (validated: mean error 0.47%). Stage 2: for a massless field with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$, both $\hbar$ and $\omega$ cancel identically, giving $\sigma=c\sqrt{\gamma}$, the massless cancellation (validated: C9 photon field, constant $\sigma$ to within 1% across seven frequencies). Stage 3: for gravity, the self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{P}$ is the unique Planck-scale rate, giving $\sigma=\ell_{P}/\sqrt{t_{P}}$.
The dimensionless prefactor $\alpha$ in $\gamma_{\mathrm{grav}}=\alpha/t_{P}$ is determined by three independent arguments (Appendix C, §C.15). Metric self-consistency requires $\langle h^{2}\rangle=\alpha^{2}\leq 1$ over one Planck time, giving the upper bound $\alpha\leq 1$. The critical damping condition: the unique frequency at which $\gamma(\omega)=t_{P}^{2}\omega^{3}$ equals $\omega$, yields $\omega^{*}=1/t_{P}$ exactly, giving $\alpha=1$ as an algebraic identity. Singularity resolution requires that curvature fluctuations reach the Planck curvature at the minimum scale, giving $\alpha\geq 1$. Together: $\alpha\leq 1\wedge\alpha=1\wedge\alpha\geq 1\Rightarrow\alpha=1$ uniquely. The complete derivation chain is developed in Appendix C.
@@ 4.4 Axiomatic Formalisation
A conceptual transition is required before stating the axioms. In §2–§3, the stochastic differential equations describe matter evolving in a fixed spacetime, and the evolution parameter is physical time measured by a laboratory clock. From this point forward, the stochastic differential equation describes spacetime itself, and the evolution parameter $\tau$ is a pre-geometric index: a label ordering the steps of a Wiener process on Wheeler’s superspace. Physical time emerges as a derived quantity, constructed from the quadratic variation of the metric process (§D.5). The reader should not identify $\tau$ with the time of any clock; clocks are built from the metric, which is the stochastic variable. This distinction is the framework’s resolution of the problem of time in canonical quantum gravity.
The empirical pattern (§3) and its physical interpretation (§§4.24.3) motivate a formal axiomatic structure. Four canonical axioms are stated:
@@
Physical time is not assumed. It emerges as the quadratic variation of the stochastic process: $t(\tau)=(1/\ell_{P}^{2}n_{\mathrm{eff}})\int\langle\mathrm{d}g,K^{-1}\mathrm{d}g\rangle$ (Appendix D, Theorem 1). By the Itô product rules, $(\mathrm{d}\tau)^{2}=0$, $\mathrm{d}\beta_{n}\mathrm{d}\tau=0$, $\mathrm{d}\beta_{m}\mathrm{d}\beta_{n}=\delta_{mn}\mathrm{d}\tau$, the drift $\mathcal{D}_{ij}$ contributes identically zero to the quadratic variation. The emergent time functional is therefore strictly monotone (from positivity of $K$), additive (from independent increments), coordinate-independent (from tensorial covariance), and exactly drift-independent (from the Itô calculus). In the macroscopic limit $L\gg\ell_{P}$, the emergent time coincides with proper time through the standard ADM relation $\mathrm{d}t_{\mathrm{proper}}=N\,\mathrm{d}t$, with stochastic corrections of $O(\ell_{P}^{2}/L^{2})\sim 10^{-70}$ for astrophysical scales. This construction addresses the problem of time in canonical quantum gravity: time is not a background coordinate but a derived quantity, constructed from the diffusion itself, with the deterministic gravitational dynamics contributing nothing to its rate.
@@ 4.5 The Superspace Diffusion Framework
The derivations that follow from axioms A1–A4 constitute the Superspace Diffusion Framework. The superspace diffusion hypothesis (Eq. 9) is one consequence, where the consequences for galactic kinematics are developed in a companion paper (Part II).
The Superspace Diffusion Framework extends the $\sigma$-continuum from an empirical observation to a formal structure derived non-parametrically from first principles.

### 25 Mar 2026 v2

@@ 3.6 Quantum Field Theory: The Roles of $\hbar$, $c$, and $G$
The pipeline was applied at seven frequencies spanning a decade ($\omega\in[10^{12},10^{13}]$ rad/s) at fixed damping $\gamma=10^{11}$ s-1. As for C8, the embedding lag was set to the autocorrelation zero-crossing ($\tau_{\mathrm{ACF}}\approx 157$ steps). The recovered $\hat{\sigma}$ was constant across all frequencies to within 1% (Fig. 9c). From 50 independent pipeline runs at the reference frequency, the recovered $\hat{\sigma}/\sigma_{\mathrm{true}}=0.998$ with bootstrap 95% CI [0.996, 0.999][0.996,\,0.999], a $0.24\%$ error (Fig. 9d).
@@
For the photon, $\gamma$ is an external parameter (cavity losses, material absorption). For gravity, however, the coupling is not external: the gravitational field is self-interacting (metric fluctuation modes carry energy, and that energy gravitates by the equivalence principle), so the self-coupling rate is uniquely determined by the remaining constant, Newton’s gravitational constant $G$, through $\gamma_{\mathrm{grav}}=1/t_{P}$, where $t_{P}=\sqrt{\hbar G/c^{5}}$ is the Planck time [26]. Substitution yields $\sigma=c\sqrt{1/t_{P}}=\ell_{P}/\sqrt{t_{P}}$, which corresponds to a proper-distance fluctuation of one Planck length per Planck time. This derivation is developed in the Discussion (§4) and proved in Appendix C (the Gravitational Diffusion Theorem).
@@ 3.7 Summary of Quantitative Results
Table 2 consolidates the principal recovered quantities, bootstrap confidence intervals, and errors for all nine domains. For each domain, both the drift and diffusion channels are reported where the physics admits a meaningful comparison.
Table 2: Summary of quantitative results from 50 independent pipeline runs per domain. $\hat{\theta}/\theta$: median normalised recovery. Bootstrap 95% CIs (10 000 pairwise resamples) serve as the primary statistical assessment. For deterministic systems, $\hat{\sigma}/\mathrm{signal}\ll 1$ confirms correct classification. For scaling exponents (C5), the multiplicative $k$-NN bias cancels in the log–log slope and the CI contains the theoretical value. Ensemble CI coverage is reported for stochastic free-run ensembles (250 realisations). Tests are ordered by the progression of complexity described in §2.4.
Test Domain Channel $\hat{\theta}/\theta$ 95% CI Error Cov.
C1 Classical mech. Drift (rate) 1.0000 $[1.0000,1.0000]$ 0.0001% (det.)
Diff. ($\hat{\sigma}/\mathrm{sig.}$) $1.2\!\times\!10^{-6}$
C4 Quantum mech. Drift ($\hat{\hbar}$$\hbar$) 1.0005 $[1.0001,1.0015]$ 0.05% (det.)
Diff. ($\hat{\sigma}/\mathrm{sig.}$) $9.5\!\times\!10^{-3}$
C6 Electromagnetism Drift ($\hat{c}/c$) 1.0000 $[1.0000,1.0000]$ $3\!\times\!10^{-6}$% (det.)
Diff. ($\hat{\sigma}/\mathrm{sig.}$) $4.3\!\times\!10^{-7}$
C2 Statistical mech. Drift ($\hat{\gamma}/\gamma$) 0.971 $[0.902,0.998]$ 2.95%
Diff. ($\hat{k}_{B}T/k_{B}T$) 0.984 $[0.979,0.989]$ 1.56% 94.5%
C3 Nuclear physics Diff. ($\hat{\sigma}/\sqrt{\mu}$) 1.059 $[1.051,1.062]$ 5.9% 95.8%
C8 Quantum field th. Diff. ($\hat{\sigma}/\sigma$) 0.997 $[0.995,1.000]$ 0.27%
C5 Chemical kinetics Diff. ($\hat{\alpha}/(-0.5)$) 1.004 $[1.000,1.007]$ 0.38% 94.0%
C7 Relativistic QM Drift ($\hat{v}_{g}/v_{g}$) 0.931 $[0.922,0.945]$ 6.9%
Diff. (slope/(1)/(-1)) 0.993 $[0.988,0.999]$ 0.72%
C9 QED (photon) Diff. ($\hat{\sigma}/\sigma$) 0.998 $[0.996,0.999]$ 0.24%
[…]
@@ 4 Discussion
The Discussion proceeds through four stages. Section 4.1 identifies the $\sigma$-continuum as an empirical observation arising from the nine-domain results. Sections 4.24.3 derive the gravitational diffusion coefficient from this observation combined with three foundational results of established physics; the complete proof chain occupies Appendix C and constitutes a theorem (the Gravitational Diffusion Theorem, Theorem C.6), not an interpretive hypothesis. Section 4.4 poses four canonical axioms motivated by the preceding derivation. Appendix D demonstrates that within these axioms, every element of the superspace SDE is uniquely determined. Table 4 summarises the epistemic status of each step.
Table 4: Derivation chain: epistemic status of each step from empirical observation through theorems to axiomatic formalisation. The column “External inputs” identifies what each step assumes beyond the preceding steps. Only four canonical axioms are postulated; all other results are derived.
Step What is established Epistemic status External inputs Location
$\sigma$-continuum Diffusion coefficients across 9 domains follow systematic pattern Empirical observation None §3
FDT master formula $\sigma^{2}=\gamma\hbar\omega/m$ for damped QHO at zero temperature Theorem
Callen–Welton App. C, Lem. C.2
Massless cancellation $\sigma=c\sqrt{\gamma}$ for massless fields; $\hbar$ and $\omega$ cancel identically Theorem
(this work)
Mass–energy equiv. App. C, Thm. C.4
Dimensional uniqueness of $\gamma_{\mathrm{grav}}$ $\gamma_{\mathrm{grav}}=\alpha/t_{P}$ is the unique rate from $\{\hbar,G,c\}$ Theorem
(this work)
None App. C, Lem. C.5
Self-bath (Mori–Zwanzig) Self-coupling follows from NZ projection of superspace FP equation Theorem
(this work)
Equivalence principle App. C, Prop. C.1
$\alpha\leq 1$ (metric consistency) Perturbation theory: $\langle h^{2}\rangle=\alpha^{2}\leq 1$ over one $t_{P}$ Theorem
(this work)
Metric positivity App. C, Thm. C.22
$\alpha=1$ (critical damping) $\gamma(\omega^{*})=\omega^{*}$ has unique solution $\omega^{*}=1/t_{P}$ Theorem
(this work)
Self-coupling rate App. C, Thm. C.23
$\alpha\geq 1$ (singularity resolution) Curvature fluctuations must reach $K_{P}$ at the Planck scale Theorem
(this work)
Curvature scaling App. C, Thm. C.24
$\sigma=\ell_{P}$ (gravitational) Combining massless cancellation with $\alpha=1$ Theorem
(this work)
Above inputs App. C, Thm. C.6
Axioms A1–A4 Formal axiomatic structure for stochastic gravity Postulate §4.4
Wiener character Lévy–Khintchine + path continuity Theorem A1 App. D, Thm. D.3
Drift uniqueness Lovelock classification + classical correspondence Theorem A3 App. D, Thm. D.12
Emergent time Quadratic variation; drift-independent by Itô calculus Theorem A2 App. D, Thm. D.10
Amplitude uniqueness Three independent arguments converge on $\sigma=\ell_{P}$ Theorem A1–A4 App. D, Thm. D.18
@@ 4.1 The $\sigma$-Continuum as Empirical Evidence for Universal Stochastic Structure
The nine-domain validation of §3 establishes that the stochastic differential equation $\mathrm{d}X=\mu(X)\,\mathrm{d}t+\sigma(X)\,\mathrm{d}W$ is not solely a mathematical device but a recurring structural element across physical domains. A single non-parametric pipeline, receiving only raw time series and sampling interval, recovers the governing equations of classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, and quantum field theory, with recovery errors quantified by bootstrap confidence intervals (Table 2) and without domain-specific input.
@@
The $\sigma$-continuum (Fig. 1) is an observation arising from these nine independent experiments. Whether the pattern extends to gravitational degrees of freedom, and what determines $\sigma$ for gravity, is addressed in the following sections.
@@ 4.2 Relativistic Structure in the Diffusion Coefficient
The pipeline recovers the $1/\gamma$ suppression to 0.72% (Fig. 8d) and the group velocity to $R^{2}=0.9999$ (Fig. 8f). The independent recovery of $v_{g}$ at each momentum mitigates a potential circularity: because $\gamma$ depends on $v$, the suppression $\sigma\propto 1/\gamma$ is validated only if the velocity is itself recovered independently, not assumed. Both channels, drift and diffusion, carry relativistic content. The speed of light $c$ enters $\sigma$ through the dispersion relation, alongside $\hbar$ and $m$.
Five structural observations follow from C6–C7 taken together. Observations 1, 4, and 5 are empirical, arising from the pipeline results. Observations 2 and 3 are mathematical consequences of theorems proved in Appendix B (Theorem B.5 and Proposition B.9).
@@
Fifth, the progression from C1 ($\sigma=0$) through C2–C5 ($\sigma>0$, non-relativistic) to C7 ($\sigma=\sigma_{0}/\gamma$, relativistic) constitutes an empirically observed hierarchy. The subsequent derivation for gravitational degrees of freedom (§4.3) continues this hierarchy through the Gravitational Diffusion Theorem (Appendix C, Theorem C.6): the fluctuation–dissipation relation, the massless cancellation, and the self-coupling rate determined by the equivalence principle.
@@ 4.3 From Quantum Fields to Gravitational $\sigma$
The Gravitational Diffusion Theorem (Appendix C, Theorem C.6). For any field variable $X$ describing gravitational degrees of freedom, the diffusion coefficient is $\sigma=X_{P}/\sqrt{t_{P}}$, where $X_{P}$ is the Planck unit of $X$. For proper-distance fluctuations: $\sigma=\ell_{P}/\sqrt{t_{P}}$.
The proof proceeds in three stages, each building on validated results.
@@
where $t_{P}=\sqrt{\hbar G/c^{5}}$ is the Planck time. The self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{P}$ is not merely the unique dimensionally consistent rate. It is derived from the Mori–Zwanzig projection of the superspace Fokker–Planck equation (Appendix C, Proposition C.1): the gravitational field acts as its own bath, with the self-coupling rate determined by the cubic vertex of the Einstein–Hilbert action. The dimensional argument identifies the correct answer; the Mori–Zwanzig projection explains why it is correct. Substitution into $\sigma=c\sqrt{\gamma}$ gives
@@
In summary, the three stages are as follows. Stage 1: the quantum harmonic oscillator (C8) establishes the fluctuation–dissipation relation $\sigma^{2}=\gamma\hbar\omega/m$ (validated: $0.27\%$ error from 50 independent runs, Table 2). Stage 2: for a massless field with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$, both $\hbar$ and $\omega$ cancel identically, giving $\sigma=c\sqrt{\gamma}$, the massless cancellation (validated: C9 photon field, $0.24\%$ error, constant $\sigma$ to within 1% across seven frequencies). Stage 3: for gravity, the self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{P}$ is the unique Planck-scale rate, giving $\sigma=\ell_{P}/\sqrt{t_{P}}$. The epistemic boundary between Stages 2 and 3 is precise: the C9 pipeline results validate the mathematical formula $\sigma=c\sqrt{\gamma}$ for an externally damped massless field. Stage 3 applies the same formula to a self-coupled field, where the damping rate is not external but is determined by the equivalence principle, the most precisely tested input in the derivation chain. The self-coupling mechanism is derived, not assumed, in Appendix C (Proposition C.1 and Theorem C.16); the pipeline validates the formula, while the appendix validates the physical mechanism.
The dimensionless prefactor $\alpha$ in $\gamma_{\mathrm{grav}}=\alpha/t_{P}$ is determined by three independent arguments (Appendix C, §C.15). Metric self-consistency requires $\langle h^{2}\rangle=\alpha^{2}\leq 1$ over one Planck time, giving the upper bound $\alpha\leq 1$. The critical damping condition: the unique frequency at which $\gamma(\omega)=t_{P}^{2}\omega^{3}$ equals $\omega$, yields $\omega^{*}=1/t_{P}$ exactly, giving $\alpha=1$ as an algebraic identity. Singularity resolution requires that curvature fluctuations reach the Planck curvature at the minimum scale, giving $\alpha\geq 1$. Together: $\alpha\leq 1\wedge\alpha=1\wedge\alpha\geq 1\Rightarrow\alpha=1$ uniquely. The three arguments share a single input, the self-coupling rate $\gamma(\omega)=t_{P}^{2}\omega^{3}$ (Theorem C.16), but are otherwise logically independent: Argument 1 invokes metric positivity, Argument 2 invokes the oscillator–diffusion boundary, and Argument 3 invokes curvature scaling. No two arguments use the same physical principle beyond the shared input. The complete derivation chain is developed in Appendix C.
@@ 4.4 Axiomatic Formalisation
The axioms require a shift in the role of the stochastic differential equation. In §2–§3, the stochastic differential equations describe matter evolving in a fixed spacetime, and the evolution parameter is physical time measured by a laboratory clock. From this point forward, the stochastic differential equation describes spacetime itself, and the evolution parameter $\tau$ is a pre-geometric index: a label ordering the steps of a Wiener process on Wheeler’s superspace. Physical time emerges as a derived quantity, constructed from the quadratic variation of the metric process (§D.5). The ordering parameter $\tau$ is not the time of any clock; clocks are built from the metric, which is the stochastic variable. This distinction is the framework’s resolution of the problem of time in canonical quantum gravity.
Appendix C derives $\sigma=\ell_{P}$ from three established results and the empirical pattern of C1–C9. The four canonical axioms below formalise this derivation into a self-contained axiomatic structure; Axiom A2 is a derived result promoted to axiomatic status for deductive economy (Appendix C, Remark following the Gravitational Diffusion Theorem).
@@
Physical time is not assumed. It emerges as the quadratic variation of the stochastic process: $t(\tau)=(1/\ell_{P}^{2}n_{\mathrm{eff}})\int\langle\mathrm{d}g,K^{-1}\mathrm{d}g\rangle$ (Appendix D, Theorem D.10). The effective mode count $n_{\mathrm{eff}}$ is formally divergent on non-compact manifolds; it is rendered finite by a Planck-scale wavelength cutoff ($n_{\mathrm{eff}}\sim V/\ell_{P}^{3}$) or by minisuperspace reduction ($n_{\mathrm{eff}}=1$ for FLRW), and is absorbed into the normalisation of $t(\tau)$ so that no physical prediction depends on $n_{\mathrm{eff}}$ separately (Appendix D, Remark D.12). By the Itô product rules, $(\mathrm{d}\tau)^{2}=0$, $\mathrm{d}\beta_{n}\mathrm{d}\tau=0$, $\mathrm{d}\beta_{m}\mathrm{d}\beta_{n}=\delta_{mn}\mathrm{d}\tau$, the drift $\mathcal{D}_{ij}$ contributes identically zero to the quadratic variation. The emergent time functional is therefore strictly monotone (from positivity of $K$), additive (from independent increments), coordinate-independent (from tensorial covariance), and exactly drift-independent (from the Itô calculus). In the macroscopic limit $L\gg\ell_{P}$, the emergent time coincides with proper time through the standard ADM relation $\mathrm{d}t_{\mathrm{proper}}=N\,\mathrm{d}t$, with stochastic corrections of $O(\ell_{P}^{2}/L^{2})\sim 10^{-70}$ for astrophysical scales. This construction addresses the problem of time in canonical quantum gravity: time is not a background coordinate but a derived quantity, constructed from the diffusion itself, with the deterministic gravitational dynamics contributing nothing to its rate.
@@ 4.5 The Superspace Diffusion Framework
The derivations that follow from axioms A1–A4 constitute the Superspace Diffusion Framework. The superspace diffusion hypothesis (Eq. 9) is one consequence. At galactic scales, the Mori–Zwanzig coarse-graining of the superspace Fokker–Planck equation (Proposition C.1) yields non-parametric predictions for the statistical distribution of gravitational acceleration; these consequences are developed in subsequent work.
The Superspace Diffusion Framework extends the $\sigma$-continuum from an empirical observation to a formal structure through a theorem chain (Appendix C) derived from the fluctuation–dissipation theorem, the equivalence principle, and the massless mode structure of linearised gravity.